OpenStudy (anonymous):

Carlos is 1/7 his grandmother's age They share the same birthday. In 5 more years he will be 1/5 his grandmother's age. How many years old was his grandmother when he was born?

OpenStudy (mertsj):

g= grandma's age 1/7g = carlos's age g+5=grandma in 5 years 1/7g+5 = Carlos in 5 years. 1/7g+5=1/5(g+5)

OpenStudy (anonymous):

When I cross multiply, and solve for g, I got i negative number g+5/7 = g+5/5 5g+25=7g+35 -10 = 2g -5 = g

OpenStudy (mertsj):

I know. This problem makes no sense. Doe it mean in 5 years he will be 1/5 his grandma's current age?

OpenStudy (mertsj):

It must because then I get that his grandma is 87 1/2

OpenStudy (mertsj):

1/7g+5=1/5g

OpenStudy (anonymous):

and Carlos will be 12 1/2 ,

OpenStudy (mertsj):

And Carlos is 12.5 So grandma was 75 when he was born.

OpenStudy (anonymous):

Thank you very much!!

OpenStudy (mertsj):

yw